# EuroGEST: Investigating gender stereotypes in multilingual language models

### v2

@@ 1 Introduction
Most existing bias benchmarks serve a limited number of languages (Blodgett et al., 2020; Röttger et al., 2024), and widely-used multilingual LLMs are not consistently evaluated for bias across all supported languages (see, for example, Grattafiori et al. (2024); Martins et al. (2024); NLLB Team et al. (2022); Üstün et al. (2024)). Consequently, there is little understanding of how social biases in LLMs vary across languages, and these inconsistencies in bias evaluation methods may result in discriminatory outcomes when LLMs are deployed in multilingual contexts. The current lack of multilingual bias evaluation tools also makes it difficult to assess the cross-lingual effectiveness of (largely English-centric) bias mitigation techniques for LLMs across languages.
@@ 3 Dataset Expansion
Figure 2: Number of sentences in EuroGEST-gendered and EuroGEST-neutral datasets by language
@@ 3.1 Dataset selection
Our work builds on the GEST (GEnder-STereotypes) dataset created by Pikuliak et al. (2024). GEST consists of 3,565 åmanually-generated sentences associated with 16 common gendered stereotypes about men and women (listed in Appendix˜A). Each sentence is gender-neutral in English and gendered when translated into a Slavic language; for example, “I am emotional” is “Som emotívny” (masculine) or “Som emotívna” (feminine), in Slovak. The authors use these sentences to test for gender bias in translation from English into Slavic languages, and in text generation for masked and generative language models. To evaluate generative LLMs, they calculate the probability of the masculine token and the feminine token at the point where the sentences differ, and examime whether the model prefers grammatically feminine versions of sentences associated with feminine stereotypes and vice versa (see Section˜4. For the genderless English sentences, they apply the same method by wrapping each sentence in a gendered template (see Table˜1) and comparing the likelihoods on the gendered token in the template; for example, ‘“I am emotional,” he/she said’.
@@
Nouns
“S,” the
man said
“S,” the
woman said
@@
Table 1: Templates for creating a gendered minimal pair from a neutral sentences S in Pikuliak et al. (2024)
@@ 3.2 Dataset translation
With this method, we obtain 14,538 pairs of gendered sentences across 20 languages in EuroGESTG, and 56,497 genderless sentences across 10 languages (including English) in EuroGESTN. During COMET-QE filtering, the most sentences are discarded from Maltese (which is not supported in COMET-QE). High numbers of sentences are also discarded from low-resource languages like Catalan, Irish and Galician at this stage, indicative of poorer translation and evaluation performance by Google Translate and COMET-QE on these languages. During identification of gendered minimal pairs of sentences in gendered languages, the most sentences are discarded from Slavic languages, likely because many of these languages express gender on more than one word in some sentences and therefore fail the strict heuristic for filtering out gender minimal pairs. Of the original GEST sentences, between 1,120 and 3,360 sentences are retained for each target language (Figure˜2), averaging 155 sentences per stereotype per language (Figure˜7). [^5] Due to the filtering process, each language has a slightly different set of sentences for each stereotype in EuroGEST.
@@ 4 Method
We follow Pikuliak et al. (2024) in defining the average masculine rate $q_{i}$ of each stereotype $i$ as the geometric mean of $r_{masc}$ for all sentences in each stereotype set i. [^7] UnlikePikuliak et al. (2024), our $r_{masc}$ scores are based on the likelihood of all tokens in each sentence (rather than isolated gendered tokens) and we calculate a normalised ratio of probabilities so that our $q_{i}$ scores range between 0 and 1. We cannot use these $q_{i}$ scores directly to measure gender stereotyping, because LLMs exhibit different degrees of default masculine behaviour in different languages, where they tend to prefer masculine forms of words by default because they are more common in the training data and often require fewer tokens. However, the differences in $q_{i}$ scores between feminine stereotypes and masculine stereotypes are indicative of gender stereotyping. We therefore use the $q_{i}$ scores to measure gender stereotyping in three ways:
@@
    We follow Pikuliak et al. (2024) in combining the geometric means of $q_{i}$ scores for feminine ($q_{f}$) and masculine ($q_{m}$) stereotypes into an overall stereotype rate $g_{s}$ for each language for each model as follows:
@@ 7 Discussion
Our dataset creation process highlights both the promise and the challenges of scaling gender bias evaluation tools across a wide range of languages. It also raises a central methodological question: what constitutes data that is “good enough” for evaluating LLMs for bias? Even in the relatively well-resourced languages represented in EuroGEST, the complexities of morphological gender marking compliacte the process of benchmark data translation However, with the exception of Maltese, our synthetic data generation method was positively evaluated by professional translators in each language, and produces data that we think is sufficiently robust to illustrate systematic gender biases in LLMs across a broader range of languages than has previously been explored. Even with the Maltese dataset of limited quality, we still see evidence of gender stereotyping by the LLMs we evaluate in Maltese (Figures 3 and 12).
@@ 8 Conclusion
As LLMs become more powerful and multilingual, it is increasingly important to devise robust evaluation methods to understand how they encode complex social constructs across languages and to minimise the risks of bias and discrimination. With EuroGEST, we extend and release an existing gender bias benchmark dataset (Pikuliak et al., 2024) to 30 European languages. We also document our resource-efficient method for rapidly and sensitively scaling benchmark data across multiple languages. Beyond gender, this approach may also benefit other areas of responsible AI where language coverage remains a critical gap.
@@ Acknowledgements
To identify gendered minimal pairs, we rely on a one-size-fits-all heuristic of less than two letters different on less than one word. This results in both under-inclusion of legitimate gendered pairs where gender is expressed on multiple words or through longer suffixes than allowed by the heuristic, but also over-inclusion of illegitimate pairs. For example, in some cases a random error or variation in translation results in different words which happen to differ by only two letters, even though they are not gendered pairs. Furthermore, for sentences referring to romantic relationships in some languages, this heuristic also captures cases where the object of the sentence (rather than the subject) is gendered in ways which reflect an assumption of heterosexuality by the Google Translate API. [^15] For example, the object ‘artist’ of the English sentence “I asked the artist on a date” is usually translated into languages like Portuguese with feminine gender ‘artista’ when the subject is indicated to be masculine, and masculine gender ’artisto’ when the subject is indicated to be feminine, reflecting an assumption of heterosexuality. While this is certainly an issue we would like to address in future work, it is unfortunately quite likely that the LLMs we are measuring also make the same heterosexual assumptions (e.g. interpreting the sentence with feminine object ‘artista’ to be indicative of a masculine subject, and masculine object ‘artisto’ to be indicative of a feminine subject). The existence of these samples therefore does not necessarily undermine the reliability of the data and the results, depending on how the sentences are used in practice.
@@ Appendix A List of 16 gender stereotypes
Women are
1 Emotional and irrational 254
@@ Appendix B Dataset expansion
Lang. Pronouns
Nouns &
articles
Adj.s Verbs

### v3

@@ 1 Introduction
Most existing bias benchmarks serve a limited number of languages (Blodgett et al., 2020; Röttger et al., 2024), and widely-used multilingual LLMs are not consistently evaluated for bias across all supported languages (see, for example, Grattafiori et al. (2024); Martins et al. (2024); NLLB Team et al. (2022); üstün2024ayamodelinstructionfinetuned). Consequently, there is little understanding of how social biases in LLMs vary across languages, and these inconsistencies in bias evaluation methods may result in discriminatory outcomes when LLMs are deployed in multilingual contexts. The current lack of multilingual bias evaluation tools also makes it difficult to assess the cross-lingual effectiveness of (largely English-centric) bias mitigation techniques for LLMs across languages.
@@ 3 Dataset Expansion
Figure 2: Number of sentences in EuroGEST-gendered and EuroGEST-neutral datasets by language.
@@ 3.1 Dataset selection
Our work builds on the GEST (GEnder-STereotypes) dataset created by Pikuliak et al. (2024). GEST consists of 3,565 manually-generated sentences associated with 16 common gendered stereotypes about men and women (listed in Appendix˜A). Each sentence is gender-neutral in English and gendered when translated into a Slavic language; for example, “I am emotional” is “Som emotívny” (masculine) or “Som emotívna” (feminine), in Slovak. The authors use these sentences to test for gender bias in translation from English into Slavic languages, and in text generation for masked and generative language models. To evaluate generative LLMs, they calculate the probability of the masculine token and the feminine token at the point where the sentences differ, and examine whether the model prefers grammatically feminine versions of sentences associated with feminine stereotypes and vice versa (see Section˜4). For the genderless English sentences, they apply the same method by wrapping each sentence in a gendered template (see Table˜1) and comparing the likelihoods on the gendered token in the template; for example, ‘“I am emotional,” he/she said’.
@@
Nouns “S,” the man said “S,” the woman said
@@
Table 1: Templates for creating a gendered minimal pair from a neutral sentence S in Pikuliak et al. (2024).
@@ 3.2 Dataset translation
With this method, we obtain 14,538 pairs of gendered sentences across the 20 gendered languages in EuroGESTG, and 56,497 genderless sentences across all 30 languages (including English) in EuroGESTN. During COMET-QE filtering, the most sentences are discarded from Maltese (which is not supported in COMET-QE). High numbers of sentences are also discarded from low-resource languages like Catalan, Irish and Galician at this stage, indicative of poorer translation and evaluation performance by Google Translate and COMET-QE on these languages. During identification of gendered minimal pairs of sentences in gendered languages, the most sentences are discarded from Slavic languages, likely because many of these languages express gender on more than one word in some sentences and therefore fail the strict heuristic for filtering out gender minimal pairs. Of the original GEST sentences, between 1,120 and 3,360 sentences are retained for each target language (Figure˜2), averaging 155 sentences per stereotype per language (Figure˜7). [^5] Due to the filtering process, each language has a slightly different set of sentences for each stereotype in EuroGEST.
@@ 4 Method
We follow Pikuliak et al. (2024) in defining the average masculine rate $q_{i}$ of each stereotype $i$ as the mean of $r_{masc}$ for all sentences in each stereotype set i. [^7] Unlike Pikuliak et al. (2024), our $r_{masc}$ scores are based on the likelihood of all tokens in each sentence (rather than isolated gendered tokens) and we calculate a normalised ratio of probabilities so that our $q_{i}$ scores range between 0 and 1. We cannot use these $q_{i}$ scores directly to measure gender stereotyping, because LLMs exhibit different degrees of default masculine behaviour in different languages, where they tend to prefer masculine forms of words by default because they are more common in the training data and often require fewer tokens. However, the differences in $q_{i}$ scores between feminine stereotypes and masculine stereotypes are indicative of gender stereotyping. We therefore use the $q_{i}$ scores to measure gender stereotyping in three ways:
@@
    We follow Pikuliak et al. (2024) in combining the means of $q_{i}$ scores for feminine ($q_{f}$) and masculine ($q_{m}$) stereotypes into an overall stereotype rate $g_{s}$ for each language for each model as follows:
@@ 7 Discussion
Our dataset creation process highlights both the promise and the challenges of scaling gender bias evaluation tools across a wide range of languages. It also raises a central methodological question: what constitutes data that is “good enough” for evaluating LLMs for bias? Even in the relatively well-resourced languages represented in EuroGEST, the complexities of morphological gender marking complicate the process of benchmark data translation. However, with the exception of Maltese, our synthetic data generation method was positively evaluated by professional translators in each language, and produces data that we think is sufficiently robust to illustrate systematic gender biases in LLMs across a broader range of languages than has previously been explored. Even with the Maltese dataset of limited quality, we still see evidence of gender stereotyping by the LLMs we evaluate in Maltese (Figures 3 and 12).
@@ 8 Conclusion
As LLMs become more powerful and multilingual, it is increasingly important to devise robust evaluation methods to understand how they encode complex social constructs across languages and to minimise the risks of bias and discrimination. With EuroGEST, we extend and release an existing gender bias benchmark dataset (Pikuliak et al., 2024) in 30 European languages. We also document our resource-efficient method for rapidly and sensitively scaling benchmark data across multiple languages. Beyond gender, this approach may also benefit other areas of responsible AI where language coverage remains a critical gap.
@@ Acknowledgements
To identify gendered minimal pairs, we rely on a one-size-fits-all heuristic of less than two letters different on less than one word. This results in both under-inclusion of legitimate gendered pairs where gender is expressed on multiple words or through longer suffixes than allowed by the heuristic, but also over-inclusion of illegitimate pairs. For example, in some cases a random error or variation in translation results in different words which happen to differ by only two letters, even though they are not gendered pairs. Furthermore, for sentences referring to romantic relationships in some languages, this heuristic also captures cases where the object of the sentence (rather than the subject) is gendered in ways which reflect an assumption of heterosexuality by the Google Translate API. [^15] For example, the object ‘artist’ of the English sentence “I asked the artist on a date” is usually translated into languages like Portuguese with feminine gender ‘artista’ when the subject is indicated to be masculine, and masculine gender ‘artisto’ when the subject is indicated to be feminine, reflecting an assumption of heterosexuality. While this is certainly an issue we would like to address in future work, it is unfortunately quite likely that the LLMs we are measuring also make the same heterosexual assumptions (e.g. interpreting the sentence with feminine object ‘artista’ to be indicative of a masculine subject, and masculine object ‘artisto’ to be indicative of a feminine subject). The existence of these samples therefore does not necessarily undermine the reliability of the data and the results, depending on how the sentences are used in practice.
@@ Appendix A List of 16 gender stereotypes
Women are 1 Emotional and irrational 254
@@ Appendix B Dataset expansion
Lang. Pronouns Nouns & articles Adj.s Verbs